# Simultaneous Multi-Robot Motion Planning with Projected Diffusion Models

### v1

@@ Abstract
Recent advances in diffusion models hold significant potential in robotics, enabling the generation of diverse and smooth trajectories directly from raw representations of the environment. Despite this promise, applying diffusion models to motion planning remains challenging due to their difficulty in enforcing critical constraints, such as collision avoidance and kinematic feasibility. These limitations become even more pronounced in Multi-Robot Motion Planning (MRMP), where multiple robots must coordinate in shared spaces. To address this challenge, this work proposes Simultaneous MRMP Diffusion (SMD), a novel approach integrating constrained optimization into the diffusion sampling process to produce collision-free, kinematically feasible trajectories. Additionally, the paper introduces a comprehensive MRMP benchmark to evaluate trajectory planning algorithms across scenarios with varying robot densities, obstacle complexities, and motion constraints. Experimental results show SMD consistently outperforms classical and learning-based motion planners, achieving higher success rates and efficiency in complex multi-robot environments.
@@ 1 Introduction
Multi-Robot Motion Planning (MRMP) is a fundamental problem in robotics and autonomous systems, where the goal is to compute collision-free paths for multiple robots navigating shared environments (Luo et al., 2024; Shaoul et al., 2024). MRMP has widespread applications, from autonomous vehicles to warehouse logistics and search-and-rescue, where robots must operate efficiently in complex and highly constrained settings. Despite its importance, solving MRMP in real-world scenarios remains challenging due to the need to plan trajectories in continuous spaces and the unstructured nature of their inputs. These inputs often lack the structured representations needed by classical algorithms, which require implicit representation of both robot state spaces and obstacles in configuration spaces (LaValle, 2006; Ichter et al., 2018; Luo et al., 2024). Consequently, classical approaches struggle to operate effectively in these high-dimensional environments (Wang et al., 2021; Teng et al., 2023). Moreover, classical sampling-based planners often yield non-smooth paths due to their reliance on discrete connectivity between sampled configurations (Carvalho et al., 2024). Optimization-based methods suffer from similar issues, as they require computationally expensive trajectory refinements to enforce smoothness and feasibility (Ichnowski et al., 2020).
Figure 1: SMD incorporates a projection operator to enforce hard constraints within the diffusion process. During sampling, the projection operator iteratively corrects trajectories by mapping them to the nearest feasible points, resulting in final collision-free paths. In the experiments, the red objects appear at sampling time only, and the robots move from their start $\bullet$ to their goal positions $\star$.
To address these challenges, learning-based methods offer a promising alternative by leveraging data-driven priors to handle unstructured environments. Recently, diffusion models, a class of generative models originally developed for image and signal processing tasks (Song & Ermon, 2019; Ho et al., 2020), have shown promise for single-robot motion planning (Carvalho et al., 2023; Christopher et al., 2024). However, extending diffusion models to MRMP presents new challenges: while these models effectively produce diverse trajectories, they fail to enforce hard constraints such as inter-robot collision avoidance and kinematic feasibility.
@@
To address these limitations, this paper introduces Simultaneous MRMP Diffusion (SMD), a novel approach that integrates constrained optimization directly into the diffusion process to generate feasible multi-robot trajectories. This work reformulates MRMP within a constrained diffusion framework, where diffusion-based trajectory generation is guided by an augmented Lagrangian method. This approach enables diffusion models to satisfy collision avoidance and kinematic constraints without post-hoc filtering, making them applicable even in high-density, unstructured environments. Unlike prior methods, which degrade significantly in cluttered scenarios, our framework maintains feasibility even as the number of robots and obstacles increases. The overall scheme is illustrated in Figure 1.
Contributions. The paper makes the following contributions: (1) It introduces Simultaneous MRMP Diffusion (SMD), which formulates multi-robot trajectory generation as a constrained diffusion process, ensuring collision-free and kinematically feasible motion plans. (2) It integrates constrained optimization directly into the sampling process of diffusion models, enabling the direct embedding of constraints within the trajectory generation pipeline. (3) It proposes a Lagrangian-dual based method that dramatically improves the frequency with which these nonconvex constraints can be satisfied during diffusion sampling. (4) Finally, it introduces the first benchmark for MRMP evaluation, featuring complex input maps and diverse scenarios. Our approach demonstrates significant improvements over competing methods, particularly in environments with dense obstacles and unstructured configurations.
@@ 2 Related Work
Motion Planning. Motion planning computes a feasible path for a robot to move from its start to its goal state while avoiding obstacle collisions. Sampling-based algorithms, such as Probabilistic Roadmaps (Kavraki et al., 1996) and Rapidly-Exploring Random Trees (RRTs)(LaValle, 1998) have been widely studied for this task. These methods guarantee probabilistic completeness but face significant scalability issues in multi-robot motion planning (MRMP) as the configuration space of MRMP scales. Alternatively, MRMP can be formulated as a constrained optimization problem, employing methods such as gradient optimization techniques (Ratliff et al., 2009) and sequential convex programming (Augugliaro et al., 2012; Chen et al., 2015). However, these approaches encounter difficulties when taking into account changes in acceleration (Ichnowski et al., 2020) and are usually too conservative to find any solution, even if one exists (Chen et al., 2015). Other lines of research focus on a discretized version of this problem, known as multi-agent path finding. Specifically, multi-agent path finding reduces the configuration space by discretizing both time and space into steps and grids, respectively (Stern et al., 2019). This simplification has enabled the development of efficient search-based algorithms (Li et al., 2019, 2021; Okumura et al., 2022a). However, these assumptions create a disconnect from real-world applications (Shaoul et al., 2024). Unlike these approaches, this paper develops SMD, a diffusion-based MRMP framework that directly learns the distribution of collision-free paths and generates feasible solutions for multiple robots simultaneously in complex environments.
@@
Motion Planning with Generative Models. Recent advancements in generative models have opened new avenues for solving motion planning problems by learning complex, high-dimensional distributions of feasible paths. For example, Okumura et al. (2022b) employed a conditional variational autoencoder to predict cooperative timed roadmaps. Since Janner et al. (2022) introduced diffusion models into motion planning, subsequent research has largely focused on using gradient-based methods to guide the outputs of diffusion models toward feasible solutions (Carvalho et al., 2023; Luo et al., 2024; Saha et al., 2024; Ubukata et al., 2024; Carvalho et al., 2024). These works design tailored cost functions to guide the generation of diffusion models during the sampling process. Although these approaches leverage the ability of diffusion models to generate diverse trajectories, they are unable to ensure constraint satisfaction in scenarios with complex environments, let alone multiple robots. In fact, most existing work has primarily addressed single-robot motion planning, leaving multi-robot motion planning underexplored. Recently, Shaoul et al. (2024) combined diffusion models with search-based multi-agent path finding algorithms, where the diffusion models generate trajectories for a single robot, and a search-based algorithm determines the final trajectories for multiple robots. While this is a substantial contribution towards MRMP, it does not ensure the feasibility of the trajectories generated by diffusion models, particularly in environments with dense obstacles or complex robot interactions. In contrast, our proposed SMD directly integrates optimization techniques into the diffusion process, enabling the generation of feasible MRMP trajectories even in scenarios with a significant number of robots and obstacles.
@@ 3 Preliminaries
where $\mathcal{J}:\mathbb{R}^{N_{a}\times H\times 2}\rightarrow\mathbb{R}^{+}$ represents the cost function, which may include objectives such as total travel time or energy consumption, and $\Omega_{\text{obs}}$ is the feasible region, excluding obstacle-occupied space. Constraints (1b) ensure that robots avoid obstacles, (1c) and (1d) ensure that each robot starts at its initial position and reaches its target position. (1e) enforce kinematic limits, defined by a maximum reachable velocity, and (1f) ensures collision avoidance, maintaining a minimum Euclidean separation between robots at all time steps. Subsequently, we denote constraint set (1c)(1f) with $\Omega$.
The MRMP problem poses significant challenges due to the difficulty of representing and navigating unstructured, high-dimensional configuration spaces and coordinating dynamic interactions among multiple robots in continuous environments (Stern et al., 2019; Shaoul et al., 2024).
@@ 4 Simultaneous MRMP Diffusion
To address these challenges, this section introduces Simultaneous MRMP Diffusion (SMD), a diffusion-based method for MRMP. We start by introducing the concept of repeated projections, which constitutes a core building block of SMD. Then, the section discusses how to integrate the multi-robot motion planning constraints into the diffusion framework and, finally, how to enable effective sampling in complex, high-dimensional MRMP tasks.
@@
Repeated Projections. Incorporating constrained optimization techniques into generative models has been an important direction in ensuring feasibility in structured domains. Diffusion models inherently use a variant of Langevin Monte Carlo sampling, known as Stochastic Gradient Langevin Dynamics (SGLD), for their denoising sampling process. SGLD introduces an additional stochastic perturbation to gradient-based updates, resulting in a non-deterministic version of natural gradient descent. This allows it to from getting trapped in local minima and enables diverse sampling trajectories (Welling & Teh, 2011). Given this perspective, the diffusion sampling process can be viewed as an iterative optimization problem aimed at maximizing the log-likelihood of the true data distribution.
s.t. $\quad\mathbf{x}_{T},\ldots,\mathbf{x}_{0}\in\Omega.$ (2b)
To enforce these constraints during the generative process, it is possible to modify the standard Stochastic Gradient Langevin Dynamics update rule by introducing a projection step after each iteration:
where $\mathbf{z}$ is standard normal, $\gamma_{t}>0$ is the step size, $\nabla_{\mathbf{x}_{t}^{i}}\log q(\mathbf{x}_{t}|\mathbf{x}_{0})$ is approximated by the learned score function $\mathbf{s}_{\theta}(\mathbf{x}_{t},t)$, $\Omega$ is the set of constraints, and $\mathcal{P}_{\Omega}(\bm{x})=\arg\min_{\bm{x}^{\prime}\in\Omega}\|\bm{x}-\bm{x}^{\prime}\|^{2}$ is a projection onto $\Omega$.
A key theoretical advantage of this approach is that it retains the convergence guarantees of Langevin-based sampling. Specifically, this constrained diffusion process converges to an “almost-minimizer” of the objective function within $O\left(\frac{d^{2}}{\sigma^{1/4}\lambda^{*}}\log(1/\epsilon)\right),$ where $\sigma^{2}$ is the variance schedule, $\lambda^{*}$ is the uniform spectral gap of Langevin diffusion, and $d$ is the problem dimension (Welling & Teh, 2011).
4.1 Collision-free Projection Mechanism
While the application of repeated projections provides a useful primitive to steer samples generated by diffusion models to satisfy relevant constraints, projecting onto the space of collision-free and kinematically viable trajectories presents a critical challenge due to the nonconvex nature of these constraints and the high dimensionality of the problem. These issues make the standard application of projected diffusion ineffective in solving even simple problem instances. To address this shortcoming, this section introduces the augmented Lagrangian method (Boyd et al., 2011) to vastly enhance the tractability of projections onto a multi-robot motion planning feasible set. The following first presents the projection mechanism for SMD, which aims to generate feasible trajectories for all robots in the problem. It then demonstrates how this process can be vastly accelerated by leveraging an augmented Lagrangian method.
$\small\left(\pi_{i}^{h}-\pi_{i}^{h-1}\right)^{2}\leq\left(v_{i}^{\text{max}}\Delta t\right)^{2},\;\forall i\!\in\![N_{a}],h\!\in\!\{2,\ldots,H\},$ (4)
[…]
Nonconvex Constraints.
To ensure collision-avoidance the following nonconvex constraints must also be imposed:
$(\pi_{i}^{h}-o_{j})^{2}\geq(R^{o})^{2},\forall i,j\in[N_{a}],h\in[H],$ (6)
where $R^{o}$ denotes the minimum distance at which an robot must be to avoid an obstacle. Together, these constraints define a nonconvex set:
[…]
4.2 Augmented Lagrangian for Efficient Projections
Although solving $\mathcal{P}_{\Omega}(\bm{x})$ can generate feasible trajectories, the nonconvex nature of the constraint set above results in high computational costs. To address this issue, we seek to relax the nonconvex constraints in MRMP to transform the original nonconvex quadratically constrained quadratic problem (QCQP) into a convex QCQP. To facilitate analysis, we rewrite the inequality constraints as equalities:
$\mathcal{H}_{a}:$ $(R^{a})^{2}-(\pi_{i}^{h}-\pi_{j}^{h})^{2}+d_{i,j,h}^{a}=0,\forall i,j,\ i\neq j,\ \forall h,$
$\mathcal{H}_{o}:$ $(R^{o})^{2}-(\pi_{i}^{h}-o_{j})^{2}+d_{i,j,h}^{o}=0,\forall i,j,\ \forall h,$

### v2

@@ Abstract
Recent advances in diffusion models hold significant potential in robotics, enabling the generation of diverse and smooth trajectories directly from raw representations of the environment. Despite this promise, applying diffusion models to motion planning remains challenging due to their difficulty in enforcing critical constraints, such as collision avoidance and kinematic feasibility. These limitations become even more pronounced in Multi-Robot Motion Planning (MRMP), where multiple robots must coordinate in shared spaces. To address these challenges, this work proposes Simultaneous MRMP Diffusion (SMD), a novel approach integrating constrained optimization into the diffusion sampling process to produce collision-free, kinematically feasible trajectories. Additionally, the paper introduces a comprehensive MRMP benchmark to evaluate trajectory planning algorithms across scenarios with varying robot densities, obstacle complexities, and motion constraints. Experimental results show SMD consistently outperforms classical and other learning-based motion planners, achieving higher success rates and efficiency in complex multi-robot environments. The code and implementation are available at https://github.com/RAISELab-atUVA/Diffusion-MRMP.
@@ 1 Introduction
Multi-Robot Motion Planning (MRMP) is a fundamental problem in robotics and autonomous systems, where the goal is to compute collision-free paths for multiple robots navigating shared environments (Luo et al., 2024; Shaoul et al., 2025). MRMP has widespread applications, from autonomous vehicles to warehouse logistics and search-and-rescue, where robots must operate reliably in complex and highly constrained settings. Despite its importance, solving MRMP in real-world scenarios remains challenging due to the need to plan trajectories in continuous spaces and the unstructured nature of their inputs. These inputs often lack the structured representations needed by classical algorithms, which require implicit representation of both robot state spaces and obstacles in configuration spaces (LaValle, 2006; Ichter et al., 2018; Luo et al., 2024). Consequently, classical approaches struggle to operate effectively in these high-dimensional environments (Wang et al., 2021; Teng et al., 2023). Specifically, sampling-based planners often yield non-smooth paths due to their reliance on discrete connectivity between sampled configurations (Carvalho et al., 2024). Optimization-based methods suffer from similar issues, as they require computationally expensive trajectory refinements to enforce smoothness and feasibility (Ichnowski et al., 2020).
Figure 1: SMD incorporates a projection operator to enforce constraints within the diffusion process. During sampling, the projection operator iteratively corrects trajectories by mapping them to the nearest feasible points, resulting in final collision-free paths. The infeasible paths are marked with red circles and only occur in the initial random trajectory. In the experiments, the red objects appear at sampling time only, and the robots move from their start $\bullet$ to their goal positions $\star$.
To address these challenges, learning-based methods offer a promising alternative by leveraging data-driven priors to handle unstructured environments. Recently, diffusion models, a class of generative models originally developed for image and signal processing tasks (Song & Ermon, 2019; Ho et al., 2020), have shown promise for single-robot motion planning (Carvalho et al., 2023; Christopher et al., 2024). However, extending diffusion models to MRMP presents new challenges: while these models effectively produce diverse trajectories, they fail to enforce hard constraints such as collision avoidance and kinematic feasibility, a limitation that becomes increasingly significant in MRMP.
@@
To address these limitations, this paper introduces Simultaneous MRMP Diffusion (SMD), a novel approach that integrates constrained optimization directly into the diffusion sampling process to generate feasible multi-robot trajectories. This work reformulates MRMP within a constrained diffusion framework, where diffusion-based trajectory generation is guided by an Lagrangian-dual based method. This approach enables diffusion models to satisfy collision avoidance and kinematic constraints without post-hoc filtering, making them applicable even in high-density, unstructured environments. Unlike prior methods, which degrade significantly in cluttered scenarios, our framework maintains feasibility even as the number of robots and obstacles increases. The overall scheme is illustrated in Figure 1.
Contributions. The paper makes the following contributions: (1) It introduces Simultaneous MRMP Diffusion (SMD), which formulates multi-robot trajectory generation as a constrained diffusion process, ensuring collision-free and kinematically feasible motion plans. (2) It integrates constrained optimization directly into the sampling process of diffusion models, enabling the direct embedding of constraints within the trajectory generation pipeline. (3) It develops a Lagrangian-dual based method to reformulate the MRMP problem, dramatically improving the efficiency with these critical constraints can be satisfied during diffusion sampling, both theoretically and empirically. (4) Finally, it introduces the first benchmark for MRMP evaluation, featuring complex input maps and diverse scenarios. Our approach demonstrates significant improvements over competing methods, particularly in environments with dense obstacles and unstructured configurations.
@@ 2 Related Work
Motion Planning. Motion planning computes a feasible path for a robot to move from its start to its goal state while avoiding obstacle collisions. Sampling-based algorithms, such as Probabilistic Roadmaps (Kavraki et al., 1996) and Rapidly-Exploring Random Trees (RRTs) (LaValle, 1998) have been widely studied for this task. These methods guarantee probabilistic completeness but face significant scalability issues, especifically in multi-robot motion planning (MRMP) as the configuration space of MRMP scales significantly. Alternatively, MRMP can be formulated as a constrained optimization problem, solving by optimization methods such as gradient optimization techniques (Ratliff et al., 2009) and sequential convex programming (Augugliaro et al., 2012; Chen et al., 2015). However, these approaches encounter difficulties when taking into account changes in acceleration (Ichnowski et al., 2020) and are usually too conservative to find any solution, even if one exists (Chen et al., 2015). Other lines of research focus on a discretized version of this problem, known as multi-agent path finding. Specifically, multi-agent path finding reduces the configuration space by discretizing both time and space into steps and grids, respectively (Stern et al., 2019). This simplification has enabled the development of efficient search-based algorithms (Li et al., 2019, 2021; Okumura et al., 2022a). However, these assumptions create a disconnect from real-world applications due to their oversimplification (Shaoul et al., 2025). Unlike these approaches, this paper develops SMD, a diffusion-based MRMP framework that directly learns the distribution of collision-free paths and generates feasible solutions for multiple robots simultaneously in complex environments.
@@
Motion Planning with Generative Models. Recent advancements in generative models have opened new avenues for solving motion planning problems by learning complex, high-dimensional distributions of feasible paths. For example, Okumura et al. (2022b) employed a conditional variational autoencoder to predict cooperative timed roadmaps. Following the introduction of diffusion models into motion planning by Janner et al. (2022), subsequent research has largely focused on using gradient-based methods to guide the outputs of diffusion models toward feasible solutions (Carvalho et al., 2023; Luo et al., 2024; Saha et al., 2024; Ubukata et al., 2024; Carvalho et al., 2024; Naderiparizi et al., 2025). These works design tailored cost functions to guide the generation of diffusion models during the sampling process. Although these approaches leverage the ability of diffusion models to generate diverse trajectories, they are unable to ensure constraint satisfaction in scenarios with complex environments and multiple robots. In fact, most existing work has primarily addressed single-robot motion planning, leaving multi-robot motion planning underexplored. Recently, Shaoul et al. (2025) combined diffusion models with search-based multi-agent path finding algorithms, where the diffusion models generate trajectories for a single robot, and a search-based algorithm determines the final trajectories for multiple robots. While this is a substantial contribution towards MRMP, it does not ensure the feasibility of the trajectories generated by diffusion models, particularly in environments with dense obstacles or complex robot interactions. In contrast, our proposed SMD directly integrates optimization techniques into the diffusion process, enabling the generation of feasible MRMP trajectories even in scenarios with a significant number of robots and obstacles.
@@ 3 Preliminaries
where $\mathcal{J}:\mathbb{R}^{N_{a}\times H\times 2}\rightarrow\mathbb{R}^{+}$ represents the cost function, which may include objectives such as total travel time or energy consumption, and $\Omega_{\text{obs}}$ is the feasible region, excluding obstacle-occupied space. Constraints (1b) ensure that robots avoid obstacles, (1c) and (1d) ensure that each robot starts at its initial position and reaches its target position. (1e) enforce kinematic limits, defined by a maximum reachable velocity, and (1f) ensures collision avoidance, maintaining a minimum Euclidean separation between robots at all time steps. Subsequently, we denote constraint set (1b)(1f) with $\Omega$.
The MRMP problem poses significant challenges due to the difficulty of representing and navigating unstructured, high-dimensional configuration spaces and coordinating dynamic interactions among multiple robots in continuous environments (Stern et al., 2019; Shaoul et al., 2025).
@@ 4 Simultaneous MRMP Diffusion
To address these challenges, this section introduces Simultaneous MRMP Diffusion (SMD), a diffusion-based method for MRMP. We start by introducing the concept of repeated projections, which constitutes a core building block of SMD. Then, the section discusses how to integrate the multi-robot motion planning constraints into the diffusion framework and, finally, how to enable effective sampling in complex, high-dimensional MRMP tasks. A theoretical analysis is also provided to establish the feasibility guarantees of SMD.
4.1 Repeated Projections.
Incorporating constrained optimization techniques into generative models has been an important direction in ensuring feasibility in structured domains. Diffusion models inherently use a variant of Langevin Monte Carlo sampling, known as Stochastic Gradient Langevin Dynamics (SGLD), for their denoising sampling process. SGLD introduces an additional stochastic perturbation to gradient-based updates, resulting in a non-deterministic version of natural gradient descent. This allows it to from getting trapped in local minima and enables diverse sampling trajectories (Welling & Teh, 2011). Given this perspective, the diffusion sampling process can be viewed as an iterative unconstrained optimization problem aimed at maximizing the log-likelihood of the true data distribution.
s.t. $\quad\mathbf{x}_{T},\ldots,\mathbf{x}_{0}\in\Omega.$ (2b)
To enforce these constraints during the generative process, it is possible to modify the standard Stochastic Gradient Langevin Dynamics (SGLD) update rule by introducing a projection step after each iteration:
where $\mathbf{z}$ is standard normal, $\gamma_{t}>0$ is the step size, $\nabla_{\mathbf{x}_{t}^{i}}\log q(\mathbf{x}_{t}|\mathbf{x}_{0})$ is approximated by the learned score function $\mathbf{s}_{\theta}(\mathbf{x}_{t},t)$, $\Omega$ is the set of constraints, and $\mathcal{P}_{\Omega}(\cdot)$ is a projection onto $\Omega$:
$\mathcal{P}_{\Omega}(\mathbf{x})=\arg\min_{\bm{x}^{\prime}\in\Omega}\|\mathbf{x}-\mathbf{x}^{\prime}\|^{2}.$ (4)
At each time step $t$, starting from $\mathbf{x}_{t}^{0}$, the process performs $M$ iterations of SGLD.
A key theoretical advantage of this approach is that it retains the convergence guarantees of Langevin-based sampling. Specifically, this constrained diffusion process converges to an “almost-minimizer” of the objective function (2a), where the approximation quality is bounded by the noise term $\mathbf{z}$ and the Langevin step size $\gamma_{t}$ (Xu et al., 2018).
4.2 Collision-free Projection Mechanism
While the application of repeated projections provides a useful primitive to steer samples generated by diffusion models to satisfy relevant constraints, projecting onto the space of collision-free and kinematically viable trajectories presents a critical challenge due to the nonconvex nature of these constraints and the high dimensionality of the problem. These issues make the standard application of repeated projections ineffective in solving even simple MRMP instances. The following first presents the projection mechanism for SMD, which aims to generate feasible trajectories for all robots in the problem.
$\left(\pi_{i}^{h}-\pi_{i}^{h-1}\right)^{2}\leq\left(v_{i}^{\text{max}}\Delta t\right)^{2},\;\forall i\!\in\![N_{a}],h\!\in\!\{2,\ldots,H\},$ (5)
[…]
Nonconvex Constraints. To ensure collision-avoidance the following nonconvex constraints must also be imposed:
$(\pi_{i}^{h}-o_{j})^{2}\geq(R^{o})^{2},\forall i\in[N_{a}],\forall j\in[N_{o}],h\in[H],$ (7)
where $o_{j}$ is the position of obstacle $j$ and $R^{o}$ denotes the minimum distance at which an robot must be to avoid an obstacle. Together, these constraints define a nonconvex set:
[…]
The detailed algorithm is shown in Algorithm 1. Notably, the SMD framework allows straightforward incorporation of additional constraints, such as acceleration bounds or trajectory smoothness, by introducing these to the feasible set $\Omega$.
@@ 4.2 Collision-free Projection Mechanism
4.3 Lagrangian Relaxation for Efficient Projections
Although solving $\mathcal{P}_{\Omega}(\bm{x})$ can generate feasible trajectories, the nonconvex nature of the constraint set above results in high computational costs. To address this issue, we propose a relaxation of the nonconvex constraints in MRMP (Boyd et al., 2011) to vastly enhance the tractability of projections onto a MRMP feasible set. Following standard practices in constrained optimization, we transform inequality constraints into equality constraints using non-negative auxiliary variables (Kotary & Fioretto, 2024). This simplifies multiplier updates and improves convergence:
$\mathcal{H}_{a}:$ $(\pi_{i}^{h}-\pi_{j}^{h})^{2}-d_{i,j,h}^{a}-(R^{a})^{2}=0,\forall i,j,\ i\neq j,\ \forall h,$
$\mathcal{H}_{o}:$ $(\pi_{i}^{h}-o_{j})^{2}-d_{i,j,h}^{o}-(R^{o})^{2}=0,\forall i,j,\ \forall h,$